a)If the sum of p terms of an A.Pis ap^2+bp, find the common difference?
b)if the sum of nterms of an A.P is 3n^2+5n. what is the common difference?

Grade:10

1 Answers

Arun
5 years ago
Dear student
 
Please ask only one question in one thread.
 
2.
 
Sn = 3n² + 5n 
Hence
 
T1 = S1 = 3+5 = 8
 
T1 + T2 = S2 = 22
 
Hence
T2 = 14
 
Hence d = 14 - 8 = 6